OpenStudy (arbyj):

algebra 2 help pleaseee im desperate

OpenStudy (arbyj):

\[y >\frac{ -1 }{ 3 } x^2 -6\] i know the vertex for this is 0,6 but how do i find another point?

OpenStudy (daniel.ohearn1):

Which point is that, which you're trying to find?

OpenStudy (arbyj):

to make the parabola

OpenStudy (daniel.ohearn1):

Even though it's a parabola, it's going to have 1/3 the slop of your typical y=x^2 parabola. Try starting with small simple numbers like 1,2,3,4...